Subject: Problem raith SNF 2005-02-27 01:17:12: EHT trip this afternoon system back up at 0114
Date: Sun, 27 Feb 2005 01:17:13 -0800

Previous user encountered EHT trip at 10:41 AM.
Computer problems resulted in communications and ELPHY + logic and several resets were
required to restore system functions and beam blanking operations.
LEO detector configuration repairs and saved to defaults and SEM detector function was
restored.(FAQ No. 3)
System up 1:00 AM user will come on at 4:00 AM and the queue will remain as it is on
CORAL with this offset.
JWC 02-27-2005:0121